# On the convergence of N-body simulations of the Solar System

### v1

@@ 6 Conclusions
In this paper we presented results from several thousand 5 Gyr integrations of the Solar System with different timesteps. Our results show that simulations with timesteps of 6.5 days or less are fully converged. For fully converged simulations, no statistically significant differences can be observed even after 5 Gyr.
@@
Surprisingly, we find that even simulations with timesteps as large as 32 days are almost indistinguishable after 5 Gyr in terms of instability fraction, time of instability, and secular frequencies. Such partially converged simulations might give unreliable results over short timescales, but the chaotic diffusion in the Solar System comes to our rescue over longer timescales. The diffusion washes out any information about the initial conditions. Any error we introduce because of the large timestep doesn’t matter as long as it remains small compared to the physical effects driving the planets’ evolution.
@@
This argument shouldn’t come as a surprise because, regardless of the timestep we use, we always introduce some amount of numerical noise ($10^{-16}$) into the simulation at every mathematical operation due to finite floating point precision. The novel result from this work is that we were able to convincingly show that numerical artifacts several order of magnitude larger than floating point precision and originating from a finite timestep do also not affect the planets’ long-term evolution.
@@
Previous studies (Wisdom, 2015; Hernandez et al., 2022) have argued that one might need up to 20 timesteps per pericenter passage time to avoid timestep resonances and achieve reliable results. This corresponds to a timestep of less than 3 days for the Solar System. Wisdom (2015) identifies potentially problematic timestep resonances, but looked only at relatively short term integrations (thousands of orbits) and very clean test cases (the Stark problem with one planet and a system with two giant planets). Whereas these test cases contain mean motion resonances, they do not contain the kind of secular resonances that are driving the long-term evolution of the Solar System (Mogavero and Laskar, 2022). The crucial difference is the timescale. For planets with masses similar to those of the (inner) Solar System, secular timescales are at least two orders of magnitude longer than those of mean motion resonances (Tamayo and Hadden, 2025). Thus any potential timestep resonance in the Solar System will be much weaker than in a system with strong mean motion resonances because the secular timescale and the timestep are so well separated. Another difference between such clean test cases and an integration of the full Solar System is the number of planets. In contrast to for example the Stark problem, the many different planet-planet interactions in the Solar System constantly change the orbits of planets and secular frequencies. Because of that the Solar System frequently enters a secular resonance only to shortly afterwards leave it again (Lithwick and Wu, 2011). So even if at one instance, a timestep resonance occurs, the system might not be in it long enough to be significantly affected by it. As this paper only looks at data from Solar System integrations, a primarily secularly evolving system, the results should not be blindly applied to other systems, especially those with strong mean motion resonances.
@@
An important point to raise here is that all integrations in this paper were performed with a symplectic integrator. Because of chaotic diffusion, the Solar System is undergoing a random walk. The variations due to this random walk reach order unity on a short timescale of only a few 100 Myr (see Fig. 3). Thus there is no need to get every individual integration exactly right. Several times during a 5 Gyr integration, we loose any information about both the initial conditions and any potential numerical error we made. One only needs to make sure that numerical artifacts (e.g. timestep resonances, finite floating point precision) do not dominate over physical effects (e.g. mean-motion resonances, secular resonances, chaotic diffusion). This is further supported by the fact that the WHCKL integrations provide identical results compared to the plain WHFast integrations, despite the fact that WHCKL more accurate in short term integrations (Rein et al., 2019). We note however that it is important that any superimposed numerical effect doesn’t lead to long-term drifts. This is the reason why symplectic integrators and unbiased implementations are crucial for this task (Brouwer, 1937; Rein and Tamayo, 2015). The conclusion from this paper is that numerical artifacts are negligible for Solar System integrations if the timestep is 6.5 days or less. Even timesteps as large as 32 days only increase the instability rate by a factor of 2 over 5 Gyr because chaotic diffusion still mostly dominates over numerical artifacts.
@@
We find that the historically often cited metric of the relative energy error can only provide a clue whether something has gone catastrophically wrong, but cannot be used to reliably determine whether physical results of simulations are converged or not. Instead, we strongly recommend that researchers always verify that the physical results they actually care about are independent of the timestep used. As we have shown in this paper, a surprisingly large timestep can result in converged results. Thus, a computationally inexpensive way to do a convergence study is to rerun a simulation with an increased timestep (which takes less time than the original simulation). If the physical results don’t change, then the simulations are likely converged. If they do change, then additional simulations with a smaller timestep should be run.
[…]
5
@@
The numerical integration methods underlying this article are part of the REBOUND package, available at https://github.com/hannorein/rebound. The code for the modified Fourier transform is available at https://boulder.swri.edu/~davidn/fmft/fmft.html. This research was made possible by the open-source projects Jupyter (Kluyver et al., 2016), iPython (Pérez and Granger, 2007), and matplotlib (Hunter, 2007; Droettboom et al., 2016).
[…]
5
@@ Acknowledgments
We thank Daniel Tamayo, Nathan Kaib, and Sean Raymond for helpful discussions and feedback on this paper. This research has been supported by the Natural Sciences and Engineering Research Council (NSERC) Discovery Grants RGPIN-2014-04553 and RGPIN-2020-04513. This research was enabled in part by support provided by Digital Research Alliance of Canada (formerly Compute Canada; alliancecan.ca). Computations were performed on the Niagara supercomputer (Loken et al., 2010; Ponce et al., 2019) at the SciNet HPC Consortium (scinethpc.ca). SciNet is funded by the following: the Canada Foundation for Innovation; the Government of Ontario; Ontario Research Fund – Research Excellence; and the University of Toronto.
[…]
5

### v2

@@ 6 Conclusions
In this paper we presented results from several thousand 5 Gyr integrations of the Solar System with different timesteps. Our results show that simulations with timesteps of 6.5 days or less are statistically converged. For these simulations, no statistically significant differences can be observed even after 5 Gyr.
@@
Surprisingly, we find that even simulations with timesteps as large as 32 days are statistically almost indistinguishable after 5 Gyr in terms of instability fraction, time of instability, and secular frequencies. Such partially statistically converged simulations might give unreliable results over short timescales, but the chaotic diffusion in the Solar System comes to our rescue over longer timescales. The diffusion washes out any information about the initial conditions. Any error we introduce because of the large timestep doesn’t matter as long as it remains small compared to the physical effects driving the planets’ evolution.
@@
This argument shouldn’t come as a surprise because, regardless of the timestep we use, we always introduce some amount of numerical noise ($10^{-16}$) into the simulation at every mathematical operation due to finite floating point precision. The novel result from this work is that we were able to show that numerical artifacts several order of magnitude larger than floating point precision and originating from a finite timestep do also not affect the planets’ long-term evolution.
Note that our goal was to determine the probability of orbits changing significantly. The specific criteria we chose (see Sect. 3) looks at whether two planets enter their mutual Hill sphere [^2] See also Table 1 of the Supplementary Information of Laskar & Gastineau (2009) which compares different instability criteria.. If our goal were to resolve actual physical collisions, our results would change significantly as none of the integrations with a fixed timestep of a few days is able to resolve such events.
@@
Previous studies (Wisdom, 2015; Hernandez et al., 2022) have argued that one might need up to 20 timesteps per pericenter passage time to avoid timestep resonances and achieve reliable results. This corresponds to a timestep of less than 3 days for the Solar System. Wisdom (2015) identifies potentially problematic timestep resonances, but looked only at relatively short term integrations (thousands of orbits) and very clean test cases (the Stark problem [^3] The Stark problem is a two body problem where one of the bodies has an additional acceleration of constant magnitude and direction, see e.g. Lantoine & Russell (2011). with one planet and a system with two giant planets). Whereas these test cases contain mean motion resonances, they do not contain the kind of secular resonances that are driving the long-term evolution of the Solar System (Mogavero & Laskar, 2022). The crucial difference is the timescale. For planets with masses similar to those of the (inner) Solar System, secular timescales are at least two orders of magnitude longer than those of mean motion resonances (Tamayo & Hadden, 2025). Thus any potential timestep resonance in the Solar System will be much weaker than in a system with strong mean motion resonances because the secular timescale and the timestep are so well separated. Another difference between such clean test cases and an integration of the full Solar System is the number of planets. For example, in contrast to the Stark problem, the many different planet-planet interactions in the Solar System constantly change the orbits of planets and secular frequencies. Because of that the Solar System frequently enters a secular resonance only to shortly afterwards leave it again (Lithwick & Wu, 2011). So even if at one instance, a timestep resonance occurs, the system might not be in the resonance long enough to be significantly affected by it. As this paper only looks at data from Solar System integrations, a primarily secularly evolving system, the results should not be blindly applied to other systems, especially those with strong mean motion resonances.
@@
An important point to raise here is that all integrations in this paper were performed with a symplectic integrator. Because of chaotic diffusion, the Solar System is undergoing a random walk. The variations due to this random walk reach order unity on a short timescale of only a few 100 Myr (see Fig. 3). Thus there is no need to get every individual integration exactly right. Several times during a 5 Gyr integration, we loose any information about both the initial conditions and any potential numerical error we made. One only needs to make sure that numerical artifacts (e.g. timestep resonances, finite floating point precision, numerical diffusion) do not dominate over physical effects (e.g. mean-motion resonances, secular resonances, chaotic diffusion). This is further supported by the fact that the WHCKL integrations provide identical results compared to the plain WHFast integrations, despite the fact that WHCKL is more accurate in short term integrations (Rein et al., 2019). We note however that it is important that any superimposed numerical effect doesn’t lead to long-term drifts (Trani et al., 2024). This is the reason why symplectic integrators and unbiased implementations are crucial for this task (Brouwer, 1937; Rein & Tamayo, 2015). The conclusion from this paper is that numerical artifacts are negligible for Solar System integrations if the timestep is 6.5 days or less. Even timesteps as large as 32 days only increase the instability rate by a factor of 2 over 5 Gyr because chaotic diffusion still mostly dominates over numerical artifacts.
@@
We find that the historically often cited metric of the relative energy error can only serve as a necessary condition to ensure that nothing has gone catastrophically wrong, but cannot be used to reliably determine whether physical results of simulations are statistically converged or not. Instead, we strongly recommend that researchers always verify that the physical results they actually care about are independent of the timestep used. This is of course also not a sufficient condition to ensure statistical convergence: all tested timesteps might be too large. However, it is another necessary condition which can provide valuable insights. And as we have shown in this paper, a surprisingly large timestep can result in converged results. Thus, a computationally inexpensive way to gather further evidence for statistically converged results is to rerun a simulation with an increased timestep (which takes less time than the original simulation). If the results change, then additional simulations with a smaller timestep should be run.
@@
The numerical integration methods underlying this article are part of the REBOUND package, available at https://github.com/hannorein/rebound. The code for the modified Fourier transform is available at https://boulder.swri.edu/~davidn/fmft/fmft.html. This research was made possible by the open-source projects Jupyter (Kluyver et al., 2016), iPython (Pérez & Granger, 2007), and matplotlib (Hunter, 2007; Droettboom et al., 2016).
@@ Acknowledgments
We thank Daniel Tamayo, Nathan Kaib, and Sean Raymond for helpful discussions and feedback on this paper. We thank two anonymous referees whose constructive feedback significantly improved the manuscript. This research has been supported by the Natural Sciences and Engineering Research Council (NSERC) Discovery Grants RGPIN-2014-04553 and RGPIN-2020-04513. This research was enabled in part by support provided by Digital Research Alliance of Canada (formerly Compute Canada; alliancecan.ca). Computations were performed on the Niagara supercomputer (Loken et al., 2010; Ponce et al., 2019) at the SciNet HPC Consortium (scinethpc.ca). SciNet is funded by the following: the Canada Foundation for Innovation; the Government of Ontario; Ontario Research Fund – Research Excellence; and the University of Toronto.